Which of the following is the domain for the equation y − 4.2 = 3.5(x − 2)?
A x ≤ −2.8
B y ≤ −2.8
C −∞ ≤ x ≤ ∞
D −∞ ≤ y ≤ ∞ Found 3 solutions by Theo, greenestamps, math_tutor2020:Answer by Theo(13342) 
 equation is y - 4.2 = 3.5 * (x - 2)
add 4.2 to both sides of the equation to get:
y = 3.5 * (x - 2) + 4.2
i think selection C.
that is minus infinity <= x <= plus infinity.
this means any real value of x can be used as input to the function and that the result will be a real value of y for each and every value of x.